9. Carnes Cosmetics stock price is $30, and it recently paid a S1.00 divideid. Ts uioiu expected to grow at a rate of 30% for the next 3 years, then grow forever at a constant rate, g; and r-9%. A t what constant rate is Carnes Cosmetics stock expected to grow after Year 3? 10. Brantley Industries invests a large sum of money in R&D. As a result, it retains and reinvests all of its earnings. In other words, Brantley does not currently pay any dividends. A major pension fund is interested in buying Brantleys stock. The pension fund manager has estimated Brantleys Free Cash Flow for the next 4 years as follows: $3Million, $6Million, $8Million and Million. After the 4th year, FCF is expected to grow at a constant rate of 3%. Brantleys required return on equity is 9%, has no debt outstanding and has 7.5Million shares outstanding a) What is the terminal value of Brantley Industries stock in year 4? b) What is the present value of Brantley Industries free cash flows during the next 4 years? c) What is Brantley Industries firm value today? d) What is an estimate of Brantley Industries price per share?
Answer #1

Question 9:

D0 = 1.00

D1 =1.00*1.30 = = 1.30

D2 = 1.30*1.30 = 1.69

D3 = 1.69*1.30 = 2.197

P3 = D3*(1+g)/(r-g)

P0 = 30, r =0.09

30 = 1.30/1.09 + 1.69/1.09^2 + 2.197/(1.09)^3 + 2.197*(1+g)/(0.09-g)

25.68841316 = 2.197*(1+g)/(0.09-g)

2.311957184 -25.68841316g = 2.197 +2.197g

0.114957184 = 27.88541316g

g = 0.0041

g =0.41225%

Expected growth rate after 2 years =0.41%

Note: We have answered one full question. Only one full question will be answered at a time